OpenStudy (anonymous):

Can you tell me if this answer is right? 57-37/37 x 100% = 20%

OpenStudy (anonymous):

That is not what I got.... 57 - 37/37 x 100% = 57 - 1 x 1.00 = 56 x 1.00 = 56

OpenStudy (mathstudent55):

Be careful with order of operations. Although it did not make a difference in this problem, it does make a difference most of the time. 57 - 37/37 x 100% = 57 - 1 x 1 = 57 - 1 = 56

OpenStudy (anonymous):

I got another answer of 54%

OpenStudy (mathstudent55):

I see, you mean (57 - 37)/37 x 100%. Are you trying to find a percent difference? If you mean [(57 - 37)/37] x 100% then you get 20/37 x 100% = 54% approx.

OpenStudy (anonymous):

ok